OpenStudy (anonymous):

solve for x:2/5+3/5x=x+5/10

OpenStudy (anonymous):

put all x into one side: x+5/10 - 3/5x = 2/5 then put all number on the other side: x-3/5x = 2/5 - 5/10 now equalize all base: 10/10x-6/10x= 4/10 - 5/10 4/10x = -1/10 4x = -1 x= -1/4

OpenStudy (anonymous):

:D thamx ive been stuck on tht forever!